Water rises up to a height h1 in a capillary tube of radius r. the mass of the water lifted in the capillary tube is M. if the radius of the capillary tube is doubled, the mass of water that will rise in the capillary tube will be
(a) M (b) 2M
(c) M/2 (d) 4M

Grade:12

2 Answers

Arun
3 years ago
Explanation:Mass of water rising in the tube of radius r is given by,
M=πr2hρNow,
Surface tension, σ=hρgr2=h'ρgr'2where h' is the height to which water rises in the second tube and r' its radius.
Since r'=2r, h'=h/2
Therefore, the mass of water in the second tube is given by,M'=πr'2h'ρ=π (2r)2h2ρ=2πr2hρ=2M
 
Vikas TU
3 years ago
Mass of water rising in the tube of radius r is given by,
M=πr2hρNow,
Surface tension, σ=hρgr2=h'ρgr'2where h' is the height to which water rises in the second tube and r' its radius.
Since r'=2r, h'=h/2
Therefore, the mass of water in the second tube is given by,M'=πr'2h'ρ=π (2r)2h2ρ=2πr2hρ=2M